Let X be a continuous random variable such that 𝑓(π‘₯) =1, 𝑓or 𝑓 0<𝑋<1.

Find the following probabilities:

  1. 𝑃(𝑋>0.5)
  2. 𝑃(0.5<𝑋<0.8)
1
Expert's answer
2022-06-01T17:56:48-0400

1.


"P(X>0.5)=1-P(X\le0.5)"

"=1-\displaystyle\int_{-\infin}^{0.5}f(x)dx=1-\displaystyle\int_{0}^{0.5}1dx"

"=1-[x]\begin{matrix}
 0.5\\
 0
\end{matrix}=0.5"

2.


"P(0.5<X<0.8)=\displaystyle\int_{0.5}^{0.8}f(x)dx"

"=\displaystyle\int_{0.5}^{0.8}1dx=[x]\begin{matrix}
 0.8\\
 0.5
\end{matrix}=0.3"